OpenStudy (anonymous):

a lady had 3 eggs she sold the first 20 for 15 cents each the next for 10 cents each, she got a total 700 cents for all. How many eggs did she have in all?

OpenStudy (anonymous):

@hartnn can you help

hartnn (hartnn):

let she have x eggs 20 for 15 cents each ----> 20*15 remaining (x-20)

hartnn (hartnn):

those remaining for 10 cents each \(\to (x-20)\times 10 \)

hartnn (hartnn):

total = 700 so \(\Large 20\times 15 + 10\times (x-20) = 700\) try to find x from here

hartnn (hartnn):

i ignored this : "a lady had 3 eggs"

OpenStudy (anonymous):

igot 22.967 hbu?

hartnn (hartnn):

how ? i am getting an integer number!

hartnn (hartnn):

ofcourse, "number of eggs" should be an integer.

hartnn (hartnn):

20*15 = 300 700-300 =400 400/10 = 40

OpenStudy (anonymous):

i got 40 at first but my uncle said it was wrong

OpenStudy (anonymous):

and this is how i did it i did 20*15 = 300 700 - 300 = 400 400/10 = 40

hartnn (hartnn):

your uncle is correct

hartnn (hartnn):

haha! thats x-20 :P x-20 = 40 add 20 on both sides

OpenStudy (anonymous):

60?

OpenStudy (anonymous):

well x = 60

hartnn (hartnn):

\(\huge \color{red}{\checkmark \quad \ddot \smile}\)

OpenStudy (anonymous):

can you show it again lol i got confused

OpenStudy (anonymous):

ooooohhh nvm nvm i get it cause they sold 40 for 10 cents and 40 plus the 20 she sold before = 60 eggs right?

hartnn (hartnn):

\(\Large 20\times 15 + 10\times (x-20) = 700 \\\Large 300 + 10\times (x-20) = 700 \\ 10\times (x-20) = 700-300 \) .... okk....yes correct :)

OpenStudy (anonymous):

thank you so very much you are the best!

hartnn (hartnn):

welcome ^_^ you too are awesome ;)